OpenStudy (anonymous):

3. Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form

OpenStudy (anonymous):

if the required line // to y = -x + 6 . it means they have the same slope. can you figure out that slope?

OpenStudy (anonymous):

we have the equation of the line is y =mx +b where m is slope of the line, b is y-intercept. you can do the same with your problem. tell me what is the slope of the given line?

OpenStudy (anonymous):

wht given line (-4,2)????

OpenStudy (anonymous):

the line paralleled to the require line, y = -x +6

OpenStudy (anonymous):



OpenStudy (anonymous):

I don't know how you can get it, but it's right.

OpenStudy (anonymous):

as long as you are ok, I'm ok, if you need more explanation, call me, if you want to stop at that result, you are ok